OpenStudy (anonymous):

Put the equation in standard form for an ellipse. 9x^2-54x+16y^2+32y=47

OpenStudy (anonymous):

\[\frac{(x-3)^2}{16}+\frac{(y+1)^2}{9}=1\]

OpenStudy (anonymous):

\[\frac{(x-3)^2}{16}+\frac{(y+1)^2}{9}=1\]

OpenStudy (anonymous):

\[\frac{(x-3)^2}{16}+\frac{(y+1)^2}{9}=1\]

OpenStudy (anonymous):

oops sorry